OpenStudy (anonymous):

Please help i am really bad at these :P 1. Find -3A+6B A= [ -3 5 -6] B= [-2 6 7] [ 9 -5 3] [2 -1 -6] A. [ 21 12 -24] [-15 9 36] B. [-3 21 60] [48 -27 36] C. [ 21 12 60] [-15 -27 36] D. [-3 21 60] [-15 9 -45] 2. solve the matrix equation X-[2 -8] = [ 4 -6] [-4 2] [ 2 -8] A. [6 -14] [-2 -6] B. [ 2 -14] [ -2 -10] C. [ 6 -14] [ 6 -10 ] D. [ 6 2] [ 6 -6]

OpenStudy (amistre64):

its just multiplication and addition ... what are you not good at?

OpenStudy (anonymous):

well with the second one im really confused on how to do it

OpenStudy (amistre64):

X - M = K X = K + M

OpenStudy (amistre64):

X - [2 -8] = [ 4 -6] [-4 2] [ 2 -8] X = [ 4 -6] + [ 2 -8] [ 2 -8] [-4 2]

OpenStudy (amistre64):

adding matrixes is just adding 'like' components ...

OpenStudy (anonymous):

so bascially you just add across like for instance 4+ 2 then -6+-8 and the so on

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

if we do some moving around for convenisnence [ 4 -6] + [ 2 -8] [ 2 -8] + [-4 2] [4 -6] + [ 2 -8] ------------ [ 2 -8] + [-4 2] -------------

OpenStudy (amistre64):

*convenience lol

OpenStudy (anonymous):

oh okay i think i get it now thank you :)

OpenStudy (amistre64):

youre welcome ... and that first one is just multipying all the elements of a matrix my the number .. 3A is just the A with all its elements multiplied by 3 .... multiply then work the addition :)

OpenStudy (amistre64):

*by the number